Problem 3-35 Using ratios to construct financial statements [LO3-2] The following information is from Harrelson Inc.s, financial statements. Sales (all credit) were $30.60 million for last year Sales to total assets Total debt to total assets Current ratio Inventory turnover Average collection period Fixed asset turnover 1.80 times 40 % 2.40 times 6 times 19 days 5 times Complete the balance sheet: (Use a 360-day year. Do not round intermediate calculations. Input your answers in millions rounded to 2 decimal places.) Answer is complete but not entirely correct S in millions $ in millions $4.16Current debt Cash Accounts receivable Inventory 1.62クLong-term debt 5.10Total debt 4.53 2.27ク 6.80ク Total current assets Fixed assets $10.88 1Equity 0.20 4.16 17.00 о Total debt and stockholders UI Total assets
Answer #1
Ans. ASSETS in millions LIABILITIES in millions
Cash 4.16 Current debt 4.53
Accounts receivables 1.62 Long - term debt 2.27
Inventory 5.1 Total debt 6.8
Total current assets 10.88
Fixed assets 6.12 Equity 10.2
Total assets 17 Total debt and stockholders' equity 17
*Calculations:
1 Sales to Total assets =   Sales / Total assets
1.80   =   30.60 / Total assets
Total assets =   30.60 / 1.80
17
2 Total debt to Total assets = Total debt / Total assets
0.40   =    Total debt / 17
Total debt   =   17 * 0.40
6.8
3 Fixed assets turnover =   Sales / Fixed assets
5   =   30.60 / Fixed assets
Fixed assets =   30.60 / 5
6.12
4 Current assets = Total assets - Fixed assets
17 - 6.12
10.88
5 Current ratio = Current assets / Current liabilities
2.40 = 10.88 / Current liabilities
Current liabilities =   10.88 / 2.40
4.53
6 Total debt = Current liabilities + Long term debt
6.8 = 4.53 + Long term debt
Long term debt = 6.8 - 4.53
2.27
7 Inventory turnover = Sales / Inventory
6 = 30.60 / inventory
Inventory = 30.60 / 6
5.1
8 Average collection period = No. of days in year / Receivables turnover
19 =   360 / Receivables turnover
Receivables turnover = 360 / 19
18.947
Receivables turnover = Sales / Receivables
18.95 = 30.60 / Receivables
Receivables   =   30.60 / 18.947
1.62
9 Current assets = Inventory + Receivables + Cash
10.88 = 5.1 + 1.62 + Cash
Cash = 10.88 - 5.1 - 1.62
4.16
10 Total debt and stockholder's equity = Total debt + Equity
17 =   6.80 + Equity
Equity = 17 - 6.80
Equity   = 10.20
*Total debt and stockholder's equity = Total assets